OpenStudy (anonymous):

Use integrals to find the area of the triangle with vertices (0,5), (-2, -2), (2,2)

OpenStudy (anonymous):

1. Calculate the three lines: \[\begin{matrix} y_1=3.5x+5 \\ y_2=x-2 \\y_3=5-1,5x \end{matrix} \] 2. Calcultate the integals: \[A=\int\limits_{-2}^{0}(y_1(x)-y_2(x))dx + \int\limits_{0}^{2} (y_3(x)-y_2(x)) dx\]